OpenStudy (anonymous):

exact value of cot^-1(cot(13pi/9))

OpenStudy (anonymous):

\[\cot ^-1(\cot (\frac{ 13\pi }{ 9}))\]

OpenStudy (anonymous):

on the unit circle

OpenStudy (anonymous):

you want the exact value?

OpenStudy (anonymous):

=\[\frac{ 1 }{ \cot(\frac{ 13\pi }{ 9 }) }\]=\[\frac{ 1 }{ \cot(\frac{ 17\pi }{ 18}+\frac{ \pi }{ 2 })}\] using the formula cot(a+b)=(cotacotb-1)/cot(a)+cot(b)= \[\frac{ \cot \frac{ \pi }{ 2 } +\cot \frac{ 17\pi }{ 18 }}{ \cot \frac{ \pi }{ 2 }\cot \frac{ 17\pi }{18 }-1}\] =\[-\cot \frac{ 17\pi }{ 18}=-\cot(\frac{ 18\pi }{ 18 }-\frac{ \pi }{ 18 })\]

OpenStudy (anonymous):

=\[-\cot(\pi-\frac{ \pi }{ 18 })=\cot(\frac{ \pi }{ 18 })\]